# Causal Panel Analysis under Parallel Trends:

### v4

@@ 4.  Data and Procedure
Finally, we compute robust confidence sets (CSs) proposed by Rambachan and Roth (2023), a partial identification approach that avoids the conditional inference problem by accounting for potential PT violations when testing the null hypothesis of no treatment effect in the post-treatment period. Specifically, we use relative magnitude (RM) restrictions, which allow for post-treatment PT violations that are at most $\bar{M}$ times the size of the maximum violation estimated in pre-treatment placebo periods. In other words, the differential trend between any two consecutive post-treatment periods are assumed to be no greater than $\bar{M}$ times the maximum difference between any two placebo periods. Using the placebo periods ensures that the benchmark pre-treatment estimates are obtained in the same way as post-treatment ATT estimates, avoiding potential asymmetry in testing and treatment effect estimation (Roth, 2024). If we first assume we can decompose each estimator for the DTE $\beta_{t}$ into the actual DTE $\tau_{t}$ and a trend component $\delta_{t}$, $\beta_{t}=\tau_{t}+\delta_{t}$, then the RM assumption requires that for all periods relative to treatment $t$,
$|\delta_{t+1}-\delta_{t}|\leq\bar{M}\cdot\max_{s\in\mathcal{P}\backslash\{0\}}\{\delta_{s+1}-\delta_{s}\}$.
in which $\mathcal{P}$ is the set of placebo periods excluding the period immediately preceding treatment adoption. In our case, we use $\mathcal{P}=\{-2,-1\}$, meaning the maximum trend is given by $\max\{\delta_{0}-\delta_{-1},\delta_{-1}-\delta_{-2}\}$. Robust CSs are uniformly valid for the partially identified treatment effects under the RM restriction. Rambachan and Roth (2023) suggest $\bar{M}=1$ as a “natural benchmark” when the number of pre-treatment periods used for testing is comparable to the number of post-treatment periods, assuming that any PT violations are no worse than the observed pretrend (p. 2653). [^14] In our case, the number of post-treatment periods typically exceeds the number of placebo periods, resulting in a laxer standard, as the maximum is non-decreasing with the number of pre-treatment periods. For each study, we compute robust CSs with $\bar{M}=0.5$ to apply a lenient benchmark. For each study, we compute the robust CSs with $\bar{M}=0.5$, erring on the lenient side. We also conduct a sensitivity analysis for each such study, obtaining the “breakdown value” of $\bar{M}$, denoted as $\tilde{M}$, at which the robust CS first includes zero. Figure 2(e) marks the robust CS for the ATT estimated using the imputation estimator in pink. In this example, the robust CS excludes zero, and the breakdown value from the sensitivity analysis is $\tilde{M}=3.0$, suggesting that the estimated coethnic mobilization effect is robust to potential PT violations under the RM restriction.
@@
Overall, the results from Grumbach and Sahn (2020) appear highly robust, regardless of the choice of point and variance estimators. The PT and no-carryover-effect assumption seem plausible. The study also has sufficient power to distinguish the ATT from zero, even under potential PT violations.
5.  Systematic Assessment
@@
We perform the replication and reanalysis procedure described above for all 49 articles in our sample. This section offers a summary of our findings, with complete results for each article available in the SM. We organize our results around two main questions: (1) Are existing empirical findings based on TWFE models robust to HTE-robust estimators? (2) Is the PT assumption plausible, and do original findings remain robust to mild PT violations informed by pretrends? We also discuss other issues observed in the replicated articles, including the presence of carryover effects and sensitivity to model specifications.
@@
To examine the impact of the weighting problem caused by HTE associated with TWFE models, we first compare the estimates obtained from the imputation estimator, FEct, for all studies to those originally reported. The dissimilarity between the original and FEct estimates proxies the extent to which the negative weighting problem is consequential in practice. We choose the imputation estimator for the reason mentioned in Section 3. Most importantly, it is applicable to all studies in our sample, including those with treatment reversals. Figure 3 plots the comparison. The horizontal axis represents the originally reported TWFE estimates, and the vertical axis represents estimates from the imputation estimator, both normalized using the same originally reported SE. If the point estimates are identical, then the corresponding point should lie exactly on the 45-degree line. Red triangles represent studies where the imputation estimates are statistically insignificant at the 5% level, based on cluster-bootstrapped SEs.
@@
However, results sometimes deviate significantly. In the top left corner of Figure 3, we plot the histogram of the ratio of imputation to TWFE estimates. Although the mean and median of the ratio are close to one, at the extremes, we observe imputation estimates as small as one-fourth or as large as more than double the TWFE estimates. The most consequential deviations occur in studies that were originally near the margins of statistical significance. Additionally, we plot the ratio of SE estimates from the imputation method to TWFE in the bottom right corner of Figure 3. The median is 1.1, meaning that in the majority of cases the SE estimate from the imputation method is at least 10% larger than the SE from TWFE. The mean is 1.23, and the distribution is right-skewed; in the extreme, the ratio was almost as high as three. Combined, drops in point estimates and increases in uncertainty lead to the third pattern: When switching from TWFE to the imputation estimator, the number of studies that are statistically insignificant at the 5% level increases from one to 12 (24%).
If we restrict our attention to eleven studies with staggered treatments, we can broaden our comparison set to include more HTE-robust estimators discussed earlier. [^15] Kogan (2021) and Magaloni, Franco-Vivanco and Melo (2020) are excluded because the original specifications include additional time trends, which are not supported by HTE-robust estimators except the imputation estimator. Figure 4 visually compares the points estimates (left panel) and $z$-scores (right panel). In the left panel, all point estimates are normalized by the same reported TWFE SE for each study. In all but three studies, the estimates from all HTE-robust estimators share the same sign, though there is a noticeable amount of variation in the estimated effect size. [^16] Although a sign change is observed in Clayton and Zetterberg (2018) when using CSDID, IW, and DIDM, these estimates are negligibly small and statistically insignificant. Similarly, the not-yet-treated version of CSDID is the opposite sign but miniscule and statistically insignificant in Bischof and Wagner (2019). The estimates from IW and CSDID are also of the opposite sign for Kuipers and Sahn (2023), and they are of a much larger magnitude. The never-treated version of CSDID is also statistically significant.
@@
Note: The above figures compare reported TWFE coefficients with estimates from various alternative estimators. In the left panel, all eight estimates for each application are normalized by the same reported SE to highlight changes resulting from the use of alternative estimators. In the right panel, the estimates are divided by their respective bootstrapped SEs. To facilitate visualization, we multiply all estimates by the sign of the reported coefficient. In both figures, black (red) circles represent estimates that would be statistically significant (insignificant) at the 5% level, assuming they were treated as $z$-scores. The normalized CSDID (never treated) estimate for Kuipers and Sahn (2023), -5.36, falls out of plotting area and is therefore not shown in the left panel. Kogan (2021) and Magaloni, Franco-Vivanco and Melo (2020) are excluded because the authors’ original TWFE specifications include unit-specific linear time trends, which are not supported by the HTE-robust estimators except the imputation estimator. Some estimates are missing because of too few never-treated units. PanelMatch is excluded because it targets a different estimand.
As in Figure 3, TWFE does not appear to be systematically upward or downward biased compared to HTE-robust estimators. Another observation that carries over is that HTE-robust estimators generally require more power to reject the null hypothesis of no effect. In five of the eleven studies, at least four HTE-robust estimates per study are statistically insignificant. The left panel shows that the changes in point estimate alone are often sufficient to render the results statistically insignificant. The comparison of $z$-scores in the right panel highlights that increased uncertainties can be substantial. Combined with earlier evidence, these findings from the staggered cases suggest that HTE bias in TWFE is empirically significant and warrants careful consideration by researchers.
It is worth noting that while only a small fraction of studies in our sample (five articles, 13.5%) employ a bootstrap procedure to estimate SEs or CIs, the widely practiced cluster-robust SE typically performs adequately. This is because the number of units (clusters) is generally large, with a median of 317. However, 10 studies have fewer than 50 units; among them, two studies that were significant at the 5% level using cluster-robust SEs fell below this threshold when using cluster-bootstrapped SEs, both of which were already marginally significant. We provide comparisons of reported, cluster-robust, and cluster-bootstrapped SEs in the SM.
Figure 5: Estimated Dynamic Treatment Effects
@@
Figure 5: Estimated Dynamic Treatment Effects (Cont.)
@@
ATT: 0.05 (0.02) Liao (2023)
ATT: -0.04 (0.01) Magaloni, Franco-Vivanco and Melo (2020)
ATT: -2.28 (0.96) Marsh (2023)
ATT: -0.01 (0.004)
@@
ATT: -2.35 (0.17) Schafer et al. (2022)
ATT: -0.02 (0.003) Schubiger (2021)
ATT: 0.05 (0.01) Schuit and Rogowski (2017)
@@
Refer to caption Refer to caption Refer to caption Note: We report the estimated ATT and corresponding bootstrap SEs (in parentheses) using the imputation estimator, FEct. We provide the $p$-values from the $F$ test on the top-left corner in each figure; rejecting the null in this test indicates potential PT violations. Note that these tests are infeasible for four cases with only one pretreatment period.
@@
Although the recent methodological literature has heavily focused on HTE, PT violations—long recognized as a potential pitfall—remain a primary concern in practice. In Figure 5, we present the event-study plot based on estimates from the imputation estimator for each study in our sample. We also report the ATT estimates and their bootstrapped SEs, as well as the $p$-values of the $F$ tests used to assess pretrends. Due to space limitations, we present the event-study plots from other estimators, as well as results from the placebo tests, robust CSs, and sensitivity analyses, in the SM.
What is concerning is that, in at least six studies (12%), the PT assumption seems highly implausible. In these studies, the DTE estimates in the pre-treatment periods deviate substantially from zero compared with those in the post-treatment period, and the $F$ test rejects the null. [^17] There are cases where the $F$ test rejects, but we do not consider them highly problematic because, with a large sample size, a confounder inconsequential to the ATT estimate can still produce a small $p$-value in the $F$ test. This is why the sensitivity analysis approach is particularly useful. For the remaining studies, the CIs of pre-treatment estimates mostly cover zero, and the $F$ test and placebo test do not reject. However, this could be simply due to a lack of statistical power. Therefore, to assess the robustness of the findings, we need additional tools that simultaneously account for both the estimated pretrend and statistical power. The sensitivity analysis proposed by Rambachan and Roth (2023) does exactly that.
@@
Note: The above figures present findings from the sensitivity analysis. The sample consists of 42 studies with more than three pre-treatment periods, allowing for such an analysis. Subfigure (a) displays the $p$-values of partially identified ATT estimates using the imputation method under restricted RM PT violations with $\bar{M}=0.5$, compared to reported $p$-values for TWFE coefficients assuming PT ($\bar{M}=0$). A square root scale is used to facilitate visualization. Black (red) triangles represent studies that are statistically significant (insignificant) at the 5% level when using the imputation method under PT. Subfigure (b) shows a histogram of $\tilde{M}$, the breakdown values of $\bar{M}$; the pink bar represents 11 studies whose ATT estimates are statistically insignificant at the 5% level when using the imputation method.
We conduct a sensitivity analysis for 42 studies with at least three pre-treatment periods, for which such a test is possible. In the vast majority of these studies, when accounting for potential PT violations—even very mild ones based on estimated pretrends in the placebo periods—we cannot conclude that the ATT is statistically significantly different from zero. In other words, the ATT estimates are not large enough to be distinguished from realistic PT violations, or from estimation and sampling errors. Figure 5(a) shows that the null of no effect is rejected at the 5% level in only 11 (26%) among the 42 studies with $\bar{M}=0.5$. Figure 5(b) displays the distribution of breakdown values $\tilde{M}$. Among the 42 studies, the median is 0.15 and the mean is 0.52. Focusing only on the 29 studies that remain statistically significant at the 5% level with the imputation estimator, the median and mean are still as low as 0.35 and 0.76, respectively. Although using HTE-robust estimators with robust CSs does not necessarily “overturn” results in most of the studies we reanalyzed, it is clear that the vast majority of studies lack the power to discern the ATT from zero under very mild PT violations of realistic magnitudes, as informed by pretrends.
@@
Our reanalysis highlights several additional issues. First, the presence of missing values is widespread. Although most methodological work presumes balanced panels without missing data, in reality, many empirical studies encounter varying degrees of data missingness. Substantial differences in results for estimators that are numerically identical in balanced panels suggest that such violations may have substantial implications. We plot patterns of treatment status for each study in the SM. Based on these plots, we also see that in some studies the pattern of missingness is either seemingly nonrandom or extremely prevalent, which weakens our confidence in the respective empirical findings.
@@
Second, we perform tests for carryover effects for all studies that feature treatment reversals. If this test is rejected, it suggests that the treatment effects persist beyond the treatment periods. Among 27 articles, 5 reject the null hypothesis of no carryover effects at 5%. Part of the concern is that imputation methods and DID extensions will use control observations from previously treated units to fit the potential outcome model or as comparisons for treated observations, and if there are carryover effects, then the comparisons will become tainted. LWX (2024) note that the presence of carryover effects for a limited number of periods is less concerning, as researchers can recode treatment to persist for some time after a unit transitions out of treatment. Despite its prevalence, we observe that carryover effects do not substantially alter the findings in most studies. Specifically, when removing two post-treatment periods for studies that reject the null of no carryover effects, the ATT estimates remain similar in magnitude. Additionally, results initially deemed statistically significant continue to maintain their statistical significance (Figure A11 in the SM). Nevertheless, we recommend that researchers make it a practice to check for potential carryover effects, considering the low cost of conducting such tests and adjustments.
@@
Finally, many studies also exhibit sensitivity to model specifications. Some studies that we exclude from our sample employ one-way FE or FE at a level different from that at which treatment is assigned. Many of these findings do not hold when we reanalyze them using a TWFE model. We should clarify that this does not imply that the original results are wrong; rather, these models implicitly operate under different identifying assumptions, and there is substantial variation in how much consideration authors give to this point. Some studies do not provide a rationale for their choice to use one-way fixed effects, while others explicitly outline the type of unobserved confounders they intend to control for. In one instance, the authors inaccurately label their specification as a DID research design. The TWFE and DID estimators are generally not equivalent, and we emphasize that this difference becomes even more pronounced when fixed effects are not assigned at the level of treatment. In such cases, a TWFE specification does not correspond to even a broadly defined DID research design.
@@
In Table 4 below, we summarize the main findings of our reanalysis. The numbers represent the proportion of studies in a given journal or across all journals that fall under the respective category (hence, a lower number is better). The first two columns relate to the ratio of the imputation estimator to the TWFE estimator, which proxies the consequences of the weighting problem due to HTE. Across all journals, this ratio is less than 0.8 in 18% of studies and less than 0.5 in 6% of studies, indicating that using an HTE-robust estimator does sometimes have a substantively significant impact and is important to use at least as a robustness check.
@@
Consequence of the weigthing problem Potential PT violations Insufficient power w/. relaxed assumptions
(1) (2) (3) (4) (5) (6) (7)
Imputation-TWFE Imputation-TWFE $F$ test Placebo test TWFE not Imputation not Imputation not reject
Journal Ratio $<0.8$ Ratio $<0.5$ reject null reject null reject null reject null null w/. $\bar{M}=0.5$
$n$ 49 49 44 44 49 49 42
APSR (16) 0.00 0.00 0.40 0.40 0.00 0.10 0.70
AJPS (10) 0.11 0.00 0.21 0.29 0.00 0.12 0.69
JOP (23) 0.30 0.13 0.30 0.15 0.04 0.39 0.79
All (49) 0.18 0.06 0.30 0.25 0.02 0.24 0.74
@@
Note: Entries are proportion of studies satisfying each set of conditions. A null is deemed being rejected if $p<0.05$. In columns (1)–(2), “ratio” refers to the ratio of the Imputation (FEct) estimate to TWFE estimate for the average effect. Five studies with a single pre-treatment period are not included in the summary statistics in columns (3)–(4). Seven studies are not included in column (7) because of too few pre-treatment periods.
Columns (3) and (4) indicate the proportion of studies with clear evidence of PT violations. Specifically, the null hypothesis is rejected at the 5% significance level for the $F$ test or the placebo test. Across all journals, around a quarter of studies reject the null hypothesis using the $F$ test, with slightly fewer rejecting when using the placebo test due to the loss in power from excluding data from the placebo periods. Note, though, that failure to reject can result from insufficient power and is not sufficient to suggest that the PT holds, but it is a modest standard that credible studies should be able to clear. [^18] On the other hand, it is also important to note that rejection does not always mean a result is invalid. We reiterate that pre-trends can reflect but are not the same as the trends in the statement of the PT. Moreover, a test can sometimes be too powerful and reject due to negligibly small pre-trends. It is thus also important to consider other evidence, such as the results from the robust CSs, that can vindicate such studies. Columns (5) and (6) display the proportion of studies in which the null hypothesis of no effect is not rejected using the TWFE and imputation estimators, respectively. When TWFE is employed, this happens in only one study (2%); however, when the imputation estimator is used, this number increases to 24%, suggesting many studies in our sample are potentially underpowered. Moreover, column (7) shows that in 74% of the studies, the robust CSs for the ATT include zero with $\bar{M}=0.5$. This further underscores the fragility of results in the existing literature when we allow even very mild PT violations.

### v5

@@ Abstract
Finally, we compute the robust confidence sets (CSs) proposed by Rambachan and Roth (2023), which account for potential PT violations when testing the null hypothesis of no post-treatment effect. Specifically, we employ the relative magnitude (RM) restriction, with two modifications to accommodate the imputation method. First, we use estimates from the placebo test to ensure that benchmark pre-treatment estimates are obtained using the same approach as post-treatment ATT estimates. This alignment prevents potential asymmetry in testing and treatment effect estimation (Roth, 2024). Second, since the imputation method does not rely on a single reference period, we explicitly incorporate the placebo estimate from the last pre-treatment period ($\hat{\delta}_{0}$) to account for deviations of post-treatment estimates from earlier reference periods. Mathematically, we decompose each DTE coefficient, $\mu_{t}$, into the true treatment effect, $\tau_{t}$, and a trend (bias) component, $\delta_{t}$, such that: $\mu_{t}=\tau_{t}+\delta_{t}$. Our modified RM restriction then requires that, for all $t\geq 0$,
$\bigl{|}\delta_{t+1}-\delta_{t}\bigr{|}\leq\bar{M}\cdot\max_{\,s\in\mathcal{P}\setminus\{0\}}\bigl{|}\delta_{s+1}-\delta_{s}\bigr{|},$ (3)
where $\mathcal{P}$ is the set of placebo periods. In our application, we set $\mathcal{P}=\{-2,-1,0\}$, so the maximum violation among placebo periods is: $\max\!\Bigl{\{}\bigl{|}\delta_{0}-\delta_{-1}\bigr{|},\;\bigl{|}\delta_{-1}-\delta_{-2}\bigr{|}\Bigr{\}}.$
When $\bar{M}=0$, the RM restriction (3) implies that $\delta_{t}=\delta_{0}$ for all $t>0$, meaning that the PT violation remains fixed at the same level as in the last pre-treatment placebo period. In this case, the robust CS obtained at $\bar{M}=0$ acts as a de-biased confidence interval, using the placebo estimate from the last pre-treatment period as the benchmark for bias. Allowing $\bar{M}>0$ permits PT violations to vary over time, but constrains trend shifts between consecutive post-treatment periods to remain within $\bar{M}$ times the largest consecutive discrepancy observed during the placebo periods.
Rambachan and Roth (2023) suggest using $\bar{M}=1$ as a “natural benchmark” when the number of placebo periods is roughly equal to the number of post-treatment periods, treating any potential PT violations as no worse than those already observed (p. 2653). [^13] In our setting, the number of post-treatment periods typically exceeds the number of placebo periods, which effectively relaxes the restriction, as the maximum discrepancy among placebo periods does not decrease with the number of post-treatment periods. In our reanalysis, we first construct 95% robust CSs for each post-treatment DTE and the ATT at $\bar{M}=0$ and $\bar{M}=0.5$. Figure 2(e) illustrates these robust CSs for the estimated ATT using the imputation method. The center of the robust CSs is smaller than the point estimate because $\hat{\delta}_{0}>0$. If the CS for $\bar{M}=0$ does not include zero, as in this case, we conduct a sensitivity analysis by varying $\bar{M}$ over a wider range to determine the “breakdown value” $\tilde{M}$, which is the smallest value of $\bar{M}$ at which the robust CS first includes zero. In Grumbach and Sahn (2020), the breakdown value is $\tilde{M}=2.5$, which means that the estimated coethnic mobilization effect remains statistically distinguishable from zero unless PT violations are more than 2.5 times the largest discrepancy observed during the placebo periods.
@@
Overall, the results from Grumbach and Sahn (2020) appear highly robust, regardless of the choice of point and variance estimators. The PT and no-carryover-effect assumptions seem plausible. The study also has sufficient power to distinguish the ATT from zero, even under potential, realistic PT violations.
Systematic Assessment
@@
We perform the replication and reanalysis procedure described above for all 49 studies in our sample. This section offers a summary of our findings, with complete results for each article available in the SM. We organize our results around two main questions: (1) Are existing empirical findings based on TWFE models robust to HTE-robust estimators? (2) Is the PT assumption plausible, and do original findings remain robust to mild PT violations informed by pretrends? We also discuss other issues observed in the replicated studies, including the presence of carryover effects and sensitivity to model specifications.
@@
To examine the impact of the weighting problem caused by HTE associated with TWFE models, we first compare the estimates obtained from the imputation estimator, FEct, for all studies to those originally reported. We choose the imputation estimator for the reason mentioned ealier. Most importantly, it is applicable to all studies in our sample, including those with treatment reversals and those with additional time trends. Figure 3 plots the comparison. The horizontal axis represents the originally reported TWFE estimates, and the vertical axis represents FEct estimates, both normalized using the same originally reported SE. If the point estimates are identical, then the corresponding point should lie exactly on the 45-degree line. Red triangles represent studies where the imputation estimates are statistically insignificant at the 5% level, based on cluster-bootstrapped SEs.
@@
However, results sometimes deviate significantly. In the top left corner of Figure 3, we plot the histogram of the ratio of imputation to TWFE estimates. Although the mean and median of the ratio are close to one, at the extremes, we observe imputation estimates as small as one-fourth or as large as more than double the TWFE estimates. The most consequential deviations occur in studies that were originally near the margins of statistical significance. Additionally, we plot the ratio of SE estimates from the imputation method to TWFE in the bottom right corner of Figure 3. The median is 1.08, meaning that in the majority of cases the SE estimate from the imputation method is at least 8% larger than the SE from TWFE. The mean is 1.25, and the distribution is right-skewed; in the extreme, the ratio was almost as high as three. Combined, drops in point estimates and increases in uncertainty lead to the third pattern: When switching from TWFE to the imputation estimator, the number of studies that are statistically insignificant at the 5% level increases from one to twelve (24%).
If we restrict our attention to eleven studies with staggered treatments, we can broaden our comparison set to include more HTE-robust estimators discussed earlier. [^14] Kogan (2021) and Magaloni, Franco-Vivanco and Melo (2020) are excluded because the original specifications include additional time trends, which are not supported by HTE-robust estimators except the imputation estimator. Figure 4 visually compares the points estimates (left panel) and $z$-scores (right panel). In the left panel, all point estimates are normalized by the same reported TWFE SE for each study. In all but three studies, the estimates from all HTE-robust estimators share the same sign, though there is a noticeable amount of variation in the estimated effect size. [^15] Although a sign change is observed in Clayton and Zetterberg (2018) when using CSDID, IW, and DID_multiple, these estimates are negligibly small and statistically insignificant. Similarly, the not-yet-treated version of CSDID is the opposite sign but miniscule and statistically insignificant in Bischof and Wagner (2019). The estimates from IW and CSDID are also of the opposite sign for Kuipers and Sahn (2023), and they are of a much larger magnitude. The never-treated version of CSDID is also statistically significant.
@@
Note: The above figures compare reported TWFE coefficients with estimates from various alternative estimators. In the left panel, all eight estimates for each application are normalized by the same reported SE to highlight changes resulting from the use of alternative estimators. In the right panel, the estimates are divided by their respective bootstrapped SEs. To facilitate visualization, we multiply all estimates by the sign of the reported coefficient. In both figures, black (red) circles represent estimates that would be statistically significant (insignificant) at the 5% level, assuming they were treated as $z$-scores. The normalized CSDID (never treated) estimate for Kuipers and Sahn (2023), -5.36, falls out of plotting area and is therefore not shown in the left panel. Kogan (2021) and Magaloni, Franco-Vivanco and Melo (2020) are excluded because the authors’ original TWFE specifications include unit-specific linear time trends, which are not supported by most HTE-robust estimators except the imputation estimator. Some estimates are missing because of too few never-treated units. PanelMatch is excluded because it targets a different estimand.
As in Figure 3, TWFE does not appear to be systematically upward or downward biased compared to HTE-robust estimators. Another observation that carries over is that HTE-robust estimators generally require more power to reject the null hypothesis of no effect. In five of the eleven studies, at least four HTE-robust estimates per study are statistically insignificant. The left panel shows that the changes in point estimate alone are often sufficient to render the results statistically insignificant. The comparison of $z$-scores in the right panel highlights that increased uncertainties can be substantial. Combined with earlier evidence, these findings from the staggered cases suggest that the HTE issue regarding TWFE is empirically significant and warrants careful consideration by researchers.
It is worth noting that while only a small fraction of studies in our sample (five studies, 13.5%) employ a bootstrap procedure to estimate SEs or CIs, the widely practiced cluster-robust SE typically performs adequately. This is because the number of units (clusters) is generally large, with a median of 317. However, 10 studies have fewer than 50 units; among them, two studies that were significant at the 5% level using cluster-robust SEs fell below this threshold when using cluster-bootstrapped SEs, both of which were already marginally significant. We provide comparisons of reported, cluster-robust, and cluster-bootstrapped SEs in the SM.
Figure 5: Dynamic Treatment Effects w/ Imputation Estimator
@@
Figure 5: Dynamic Treatment Effects w/ Imputation Estimator (Cont.)
@@
ATT: 0.05 (0.02) Latura and Weeks (2023)
ATT: 0.03 (0.01) Liao (2023)
ATT: -0.045 (0.006) Magaloni, Franco-Vivanco and Melo (2020)
ATT: -2.28 (0.96)
@@
ATT: -2.16 (0.16) Schafer et al. (2022)
ATT: -0.023 (0.003) Schubiger (2021)
ATT: 0.047 (0.009) Schuit and Rogowski (2017)
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Note: We report the estimated ATT and corresponding bootstrap SEs (in parentheses) using FEct. For Skorge (2023), we use an “exit” plot becuase all treated units receives the treatment in the first period.
@@
Although the recent methodological literature has heavily focused on HTE, PT violations—long recognized as a potential pitfall—remain a primary concern in practice. In Figure 5, we present the event-study plot based on estimates from the imputation estimator for each study in our sample. We also report the ATT estimates and their bootstrapped SEs. Due to space limitations, we present the event-study plots from other estimators, as well as results from the placebo tests, robust CSs, and sensitivity analyses, in the SM.
What is concerning is that, in at least six studies (12%), the PT assumption seems highly implausible. In these studies, the DTE estimates in the pre-treatment periods deviate substantially from zero compared with those in the post-treatment period, and the $F$ test rejects the null. [^16] There are other cases where the $F$ test rejects, but we do not consider them highly problematic because, with a large sample size, a confounder inconsequential to the ATT estimate can still produce a small $p$-value in the $F$ test. This is why the sensitivity analysis approach is particularly useful. For the remaining studies, the CIs of pre-treatment estimates mostly cover zero, and the $F$ test and placebo test do not reject. However, this could be simply due to a lack of statistical power. Therefore, to assess the robustness of the findings, we need additional tools that simultaneously account for both the estimated pretrend and statistical power. The sensitivity analysis proposed by Rambachan and Roth (2023) addresses this issue. Therefore, we conduct such an analysis with a modified RM restriction (3) for 42 studies that have at least three pre-treatment periods and are suitable for this test. [^17] For studies with only three pre-treatment periods, we set the number of placebo periods in their placebo tests to 2.
@@
Note: The above figures present findings from the sensitivity analysis. The sample consists of 42 studies with more than three pre-treatment periods, allowing for such an analysis. Subfigure (a) displays the $p$-values of partially identified ATT estimates using the imputation method under restricted RM PT violations with $\bar{M}=0.5$, compared to reported $p$-values for TWFE coefficients assuming PT. A square root scale is used to facilitate visualization. Black (red) triangles represent studies that are statistically significant (insignificant) at the 5% level when using the imputation method under PT. Subfigure (b) shows a histogram of $\tilde{M}$, the breakdown values of $\bar{M}$; the dark pink bar represents studies whose ATT estimates are statistically insignificant at the 5% level when using the imputation method.
Figure 6(a) shows that when we construct a robust CS with $\bar{M}=0.5$, the null hypothesis of no effect is rejected at the 5% level in only eight (19%) of the 42 studies. Many studies that were robust to HTE now appear underpowered. Figure 6(b) displays the distribution of breakdown values $\tilde{M}$. The spike at 00 consists of two types of studies: those that are statistically insignificant with the imputation estimator when three placebo periods are considered (dark gray) and those that become statistically insignificant due to debiasing using the placebo estimate $\hat{\delta}_{0}$. In other words, many results are not robust when we take into account $\hat{\delta}_{0}$ even without a trend shift. Among the 42 studies, the median is close to 0 and the mean is 0.33. Focusing only on the studies that remain statistically significant with the imputation estimator, the median and mean are still as low as 0.10 and 0.47, respectively.
These results suggest that although most statistically significant findings are robust to the imputation estimator, in the vast majority of these studies, accounting for potential PT violations—even very mild ones based on estimated pretrends during the placebo periods—prevents us from rejecting the null. In other words, in most studies, the ATT estimates are not substantial enough to be differentiated from realistic PT violations or from estimation and sampling errors.
@@
Our reanalysis highlights several additional issues. First, the presence of missing values is widespread. Although most methodological work presumes balanced panels without missing data, in reality, many empirical studies encounter varying degrees of data missingness. Substantial differences in results for estimators that are numerically identical in balanced panels suggest that such violations may have substantial implications. During replication, we generate plots that display the patterns of treatment status for each study. Based on these plots, we also see that in some studies the pattern of missingness is either seemingly nonrandom or extremely prevalent, which weakens our confidence in the respective empirical findings.
@@
Second, we perform tests for carryover effects for all studies that feature treatment reversals. If this test is rejected, it suggests that the treatment effects persist beyond the treatment periods. Among 27 studies, 5 reject the null hypothesis of no carryover effects at 5%. Part of the concern is that imputation methods and DID extensions will use control observations from previously treated units to fit the potential outcome model or as comparisons for treated observations, and if there are carryover effects, then the comparisons will become tainted. LWX (2024) note that the presence of carryover effects for a limited number of periods is less concerning, as researchers can recode treatment to persist for some time after a unit transitions out of treatment. Despite its prevalence, we observe that carryover effects do not substantially alter the findings in most studies. Specifically, in six studies that reject the null of no carryover effects, when we exclude two periods after the treatment switches back to zero, the ATT estimates remain similar in magnitude, and statistically significant results remain significant (Figure A8 in the SM). Nevertheless, we recommend that researchers make it a practice to check for potential carryover effects, considering the low cost of conducting such tests and adjustments.
@@
Finally, many studies also exhibit sensitivity to model specifications. Some studies that we exclude from our sample employ one-way fixed effects or fixed effects at a level different from that at which treatment is assigned. Many of these findings do not hold when we reanalyze them using a TWFE model. We should clarify that this does not imply that the original results are wrong; rather, these models implicitly operate under different identifying assumptions, and there is substantial variation in how much consideration authors give to this point. Some studies do not provide a rationale for their choice to use one-way fixed effects, while others explicitly outline the type of unobserved confounders they intend to control for. In one instance, the authors inaccurately label their specification as a DID research design. The TWFE and DID estimators are generally not equivalent, and we emphasize that this difference becomes even more pronounced when fixed effects are not assigned at the level of treatment. In such cases, a TWFE specification does not correspond to even a broadly defined DID research design.
@@
In Table 4 below, we summarize the main findings of our reanalysis. The numbers represent the proportion of studies in a given journal or across all journals that fall under the respective category (hence, a lower number is better). The first two rows relate to the ratio of the imputation estimator to the TWFE estimator, which proxies the consequences of the weighting problem due to HTE. Across all journals, this ratio is less than 0.8 in 18% of studies and less than 0.5 in 6% of studies, indicating that using an HTE-robust estimator does sometimes have a substantively significant impact and is important to use at least as a robustness check.
@@
APSR AJPS JOP All $n$
Imputation to TWFE Ratio $<0.8$ 0.00 0.13 0.30 0.18 49
Imputation to TWFE Ratio $<0.5$ 0.00 0.00 0.13 0.06 49
$F$ test reject null 0.30 0.21 0.30 0.27 44
Placebo test reject null 0.40 0.29 0.15 0.25 44
TWFE not reject null 0.00 0.00 0.04 0.02 49
Imputation not reject null 0.00 0.13 0.44 0.25 49
Imputation not reject null w/ $\bar{M}=0$ 0.40 0.23 0.74 0.50 42
Imputation not reject null w/ $\bar{M}=0.5$ 0.70 0.77 0.90 0.81 42
@@
Note: Entries (except in the last column) are proportions of studies satisfying each set of conditions. A null is deemed rejected if $p<0.05$. Five studies with a single pre-treatment period are not included in the summary statistics for the $F$ test and placebo test. Seven studies are not included in the last two rows due to too few pre-treatment periods.
The third and fourth rows indicate the proportion of studies with evidence of PT violations based on the $F$ test and placebo tests, respectively. Across all journals, around a quarter of studies reject the null hypothesis at the 5% level using the $F$ test, with slightly fewer rejecting when using the placebo test due to the loss in power from excluding data from the placebo periods. Note, though, that failure to reject can result from insufficient power and is not sufficient to suggest that the PT holds.
The fifth and sixth rows display the proportion of studies in which the null hypothesis of no effect is not rejected using the TWFE and imputation estimators, respectively. When TWFE is employed, this occurs in only one study (2%); however, when the imputation estimator is used, this number increases to 24%, suggesting that many studies in our sample are potentially underpowered with an HTE-robust estimator. The last two rows show that in 50% and 74% of the studies, the robust CSs for the ATT include zero with $\bar{M}=0$ (debiasing only) and $\bar{M}=0.5$ (debiasing plus restricted trend shifts), respectively. The comparison of results in the last four rows highlights that the main source of fragility in the existing literature is potential PT violations, rather than concerns of HTE.